A football kicker can give the ball an initial spe
Answer #1

Solution:

a) 19.1 degrees

b) 75 degrees

Initial speed = v = 31 m/s

Range = X = 48 m

Height = y = 3.44 m

The equation of the trajectory is y = X tan \theta - 1/2 g X^2 / ( v cos \theta)^2

(which is derived from X = (v cos \theta) T and y = (v sin\theta )T - 1/2 g T^2)

y = X tan \theta - 1/2 g X^2 / ( v cos \theta)^2 can be written as

y =  X tan \theta - 1/2 g X^2 / ( v )^2 * (1 / cos \theta)^2

Now in the above equation, trigonometric identity 1/ cos^2 \theta= 1+tan^2 \theta is substituted.

=> y = X tan \theta - 1/2 g X^2 / ( v )^2 [(1+tan^2 \theta) ]

=> y = 48 tan \theta - 4.9 (48)^2 /(31)^2 (1 + tan^2 \theta)   {here 1/2 g = 9.8 /2 = 4.9)

=> 3.44 = 48 tan \theta - 11.74776 - 11.74776 tan^2 \theta

=> 11.74776 tan^2 \theta - 48 tan \theta + 15.18776 =0

solving this quadratic equation for tan \theta,

tan \theta = 3.74 and 0.34565

=>a) Least angle = tan ^-1( 0.34565) = 19.1 degrees

b) Maximum angle = tan^-1(3.74) = 75 degrees